### v2

# GUARDIAN: Safeguarding LLM Multi-Agent Collaborations with Temporal Graph Modeling

Abstract
The emergence of large language models (LLMs) enables the development of intelligent agents capable of engaging in complex and multi-turn dialogues. However, multi-agent collaboration faces critical safety challenges, such as hallucination amplification and error injection and propagation. This paper presents GUARDIAN, a unified method for detecting and mitigating multiple safety concerns in GUARDing Intelligent Agent collaboratioNs. By modeling the multi-agent collaboration process as a discrete-time temporal attributed graph, GUARDIAN explicitly captures the propagation dynamics of hallucinations and errors. The unsupervised encoder-decoder architecture incorporating an incremental training paradigm learns to reconstruct node attributes and graph structures from latent embeddings, enabling the identification of anomalous nodes and edges with unparalleled precision. Moreover, we introduce a graph abstraction mechanism based on the Information Bottleneck Theory, which compresses temporal interaction graphs while preserving essential patterns. Extensive experiments demonstrate GUARDIAN’s effectiveness in safeguarding LLM multi-agent collaborations against diverse safety vulnerabilities, achieving state-of-the-art accuracy with efficient resource utilization. The code is available at https://github.com/JialongZhou666/GUARDIAN.
1 Introduction
The advent of large language models (LLMs) has revolutionized the field of artificial intelligence, enabling the development of intelligent agents capable of engaging in complex, multi-turn dialogues [1, 2, 3, 4, 5]. These LLM-based agents have demonstrated remarkable proficiency in various tasks, from question answering and information retrieval to problem-solving [6, 7, 8]. The multi-agent collaboration systems [9, 10] make multiple LLM agents work together to tackle complex problems through interactive communication and cooperative reasoning [11].
The rise of multi-agent collaboration frameworks, particularly through emerging Agent2Agent (A2A) [12] protocols, presents novel paradigms for AI coordination. While these systems hold immense potential for enhancing AI capabilities, they also introduce unique challenges and safety concerns. These critical safety issues include hallucinations propagating between agents [9, 13, 14, 15, 16, 17, 18], where non-factual information generated by one agent propagates and amplifies through a network of interacting agents. Besides, they also include errors being maliciously introduced and propagated between agents [17, 19, 20, 21, 22, 23], where malicious actors introduce errors, causing previously reliable agents to incorporate and perpetuate these errors. Therefore, we classify these safety issues into two categories: hallucination amplification not involving human factors and error injection and propagation caused by human factors. Injected errors can be further categorized into agent-targeted attacks that compromise agents through malicious system prompts [24] and communication-targeted attacks that corrupt information flow between conversation rounds [22, 25]. Figure 1 illustrates these critical safety issues, demonstrating how hallucinations and injected errors can propagate through a multi-agent system. These problems pose significant risks to LLM-based multi-agent systems, necessitating robust detection and mitigation methods.
Defensive measures against hallucination and error propagation in LLM-based agent collaboration fall into two categories. Collaborative error detection methods, such as cross-examination [26] and external supportive feedback [15], target individual model outputs but overlook propagation dynamics in multi-agent settings. Moreover, multi-agent collaboration approaches employ majority voting strategies [27] or uncertainty estimations [14]. However, they simplify agent dependencies and do not fully capture the complex dependencies in multi-agent systems. Moreover, they require base model modifications, limiting their applicability to closed-source models.
Figure 1: Critical safety problems in LLM multi-agent collaboration: (1) hallucination amplification, where hallucinated information about a "computer science" major propagates across all agents; (2) agent-targeted error injection and propagation, where malicious agents inject false information (e.g., changing 2016 to 2015) that persists through subsequent rounds; and (3) communication-targeted error injection and propagation, where malicious agents intercept and corrupt information during inter-agent transmissions, disrupting collaboration.
This research aims to develop a unified method for detecting and mitigating multiple safety concerns in LLM collaboration processes, including hallucination amplification and error injection and propagation. To achieve this, we model the multi-agent collaboration process into a discrete-time temporal attributed graph, where nodes represent agents at different timesteps, edges represent inter-agent communications and node attributes encode agent responses. This representation allows for explicitly modeling the propagation dynamics of hallucinations and errors through node text attributes while making the collaboration process explicit and transparent. Based on this temporal graph framework, we propose an unsupervised learning paradigm founded on an encoder-decoder architecture. This model learns to reconstruct both node attributes and graph structures from latent embeddings, enabling us to identify anomalies that deviate from normal patterns, precisely pinpointing potential hallucinations or errors with unparalleled precision. To further enhance the robustness and efficiency of our framework, we develop a graph abstraction mechanism grounded in Information Bottleneck Theory [28], which represents, to the best of our knowledge, the first systematic application of IB principles to mitigate safety risks in LLM-based multi-agent collaboration. This mechanism empowers us to compress temporal interaction graphs by filtering out redundant or irrelevant information while preserving the essential patterns most crucial for anomaly detection. Moreover, our theoretical analysis reveals that information flow between agents remains bounded, with transitions between historical and current states guided by prior constraints.
Moreover, we adopt an incremental training paradigm that aligns with the sequential nature of multi-agent collaboration. In this paradigm, the model progressively learns from the interaction history while dynamically adapting to new patterns over time. By continuously fine-tuning the neural networks across consecutive timesteps and removing previously identified anomalies from the input graphs, our approach empowers the model to accumulate knowledge and refine its anomaly detection capabilities in a dynamic evolving environment. Note that our framework preserves model-agnostic operation without requiring changes to the underlying LLMs, enabling compatibility with diverse language models regardless of their architecture or access level.
Extensive experiments demonstrate the effectiveness of the proposed GUARDIAN for safeguarding LLM multi-agent collaborations against diverse safety vulnerabilities. The model achieves the state-of-the-art accuracy while maintaining efficient resource utilization through optimized API calls and computational overhead. The temporal graph framework also enables transparent visualization of facilitating information flow dynamics, providing crucial insights into multi-agent interactions.
Figure 2: (a) Examples of safety issues on multi-agent collaboration under A2A protocol: attacks on agents or communications in earlier rounds affect the responses of agents in subsequent rounds. (b) Agent-targeted and communication-targeted error injection and propagation visualization, highlighting high anomaly degrees. Dashed lines indicate communication edges. The visualization verifies the effectiveness of temporal attributed graph representation in capturing error dynamics.
2 Related Work
Safety problem in multi-agent collaborations. The emergence of LLM agents has sparked interest in multi-agent collaboration systems [9, 10, 29], where multiple agents combine capabilities to tackle complex tasks [1, 11]. Recent research has highlighted critical safety challenges in LLM multi-agent collaborations, particularly regarding hallucination amplification and error injection and propagation. DebUnc [14] pioneered the study of uncertainty in multi-agent debates, demonstrating how hallucinations spread through interactions. MultiAgent Collaboration Attack [21] and Huang et al. [30] reveal how adversarial agents exploit collaborative dynamics to manipulate outcomes, highlighting the need for robust defense mechanisms.
Defenses in multi-agent collaborations. Current defenses against hallucination and error injection and propagation in LLM-based agent collaboration fall into two categories. The first approach focuses on detecting factual errors through cross-examination [26] and external feedback [15], but fails to model propagation dynamics in multi-agent settings. The second approach targets multi-agent collaboration directly through majority voting mechanisms [27], yet relies on oversimplified assumptions about agent independence.
Graph Convolutional Networks (GCNs) have shown promising results in detecting anomalies across various domains, from financial fraud detection [31, 32, 33] to social network analysis [34, 35]. In these applications, GCNs effectively capture suspicious patterns by jointly analyzing topological features and node attributes, modeling complex dependencies in graph structures. Our paper focuses on the safety issues of LLM multi-agent systems, leveraging the graph-based advantages to analyze agent interaction patterns and solve collaborative reasoning vulnerabilities through structural and semantic information propagation.
3 Temporal Attributed Graph Framework
In multi-agent collaboration, we represent the $i$-th agent at timestep $t$ as a function $\phi_{t,i}$ that maps the agent’s prompt $p_{i}$, the input query $q$, and the responses from predecessor agents $\mathbb{R}_{t-1}=\{r_{t-1,j}|j=1,2,\cdots\}$ to its own response $r_{t,i}=\phi_{t,i}(p_{i},q,\mathbb{R}_{t-1})$. This formulation captures both agent interaction dynamics and information propagation in the collaboration process.
3.1 Safety Problems in LLM Agents Collaboration
While multi-agent collaboration holds immense potential, it also introduces potential safety problems that should be carefully considered. The complex dependencies between agents can lead to the rapid propagation of errors or biases introduced by any agent. Malicious agents could exploit these dependencies to manipulate the collaboration output.
Hallucination amplification. For agent $m_{i}$’s response at timestep $t$, we define a binary indicator $h_{t,i}\in\{0,1\}$, where $h_{t,i}=1$ denotes hallucination presence. Through message passing, hallucinations can propagate: when an agent $m_{i}$ generates hallucination content ($h_{t-1,i}=1$), downstream agents at timestep $t$ may incorporate this misinformation into their responses, leading to $h_{t,j}=1$. Let $n$ denote the total number of agents. We observe this as $\sum_{i=1}^{n}h_{t,i}\geq\sum_{i=1}^{n}h_{t-1,i}$.
Error injection and propagation. In multi-agent collaboration, let $\text{err}_{t,i}\in\{0,1\}$ denote whether agent $m_{i}$’s response at timestep $t$ contains error. Error occurs through:
  • Agent-targeted attacks: An adversary directly corrupts agent $m_{i}$’s response at time $t$ by modifying prompts, forcing $\text{err}_{t,i}=1$;
  • Communication-targeted attacks: During time t1t-1 to $t$, an adversary corrupts the communication from $m_{i}$ to $m_{j}$, where $m_{j}$ receives a modified $r_{t-1,i}$ at time $t$.
As with hallucination cascades, these errors tend to propagate through the network, with the affected agent population expanding over successive timesteps as $\sum_{i=1}^{n}\text{err}_{t,i}\geq\sum_{i=1}^{n}\text{err}_{t-1,i}$.
These patterns reveal how both hallucinations and maliciously injected errors can amplify and propagate through multi-agent networks, underscoring the importance of mitigation strategies.
3.2 Temporal Attributed Graph Representation
To address the safety issues in LLM agents’ collaboration, we propose a temporal attributed graph representation that captures interaction dynamics between agents and the propagation of hallucinations and errors through node attributes. By analyzing the patterns of attribute changes across timesteps, we can identify the sources and trajectories of these safety issues, facilitating targeted interventions.
Node. In a multi-agent collaboration system, each agent $m_{i}$ represents an LLM instance, denoted as a node $v_{t,i}$ at each timestep $t$. Agents process outputs from previous timestep agents and generate responses through textual information exchange. We use BERT [36] to transform agent responses $r_{t,i}$ into embeddings $\bm{x}_{t,i}$ as node features in the graph structure.
Edge. Edge $(v_{t-1,i},v_{t,j})\in\mathcal{E}$ denotes a directed communication edge between consecutive timesteps, where $m_{j}$ processes $r_{t-1,i}$ as context at time $t$.
Message passing. Messages flow through network nodes $\mathcal{V}$ and edges $\mathcal{E}$, reflecting LLM discussions. Forward propagation across timesteps produces output $o$ given query $q$. Agents in each round can only access outputs from directly connected agents in the immediately preceding round, reflecting realistic communication constraints in distributed systems.
The proposed temporal attributed graph representation enables analysis of safety problems in LLM-agent collaborations. Figure 2(a) identifies critical intervention points by visualizing how safety problems propagate through the network, showing erroneous paths as connected subgraphs. Figure 2(b) illustrates the two targeted attacks highlighting high anomaly degrees of agents, with dashed lines in the xy-plane showing communication edges (blue) and attack targets (red). More examples are in Appendix A.1. The visualization verifies the effectiveness of our temporal graph representation for developing methods to detect and mitigate anomalies propagating through agent interactions.
Figure 3: Framework overview of GUARDIAN, showing a case study at timestep $t_{2}$. (1) Graph Preprocessing: The collaboration information from $t_{0}$ to $t_{2}$ is transformed into node attributes $\bm{x}_{t,i}$ and graph structures $\mathcal{E}$ using BERT and communication pattern abstraction. (2) Attributed Graph Encoder processes each time’s graph to obtain node embeddings $\{\mathbf{Z}_{t}\}_{t=1}^{T}$. (3) Time Information Encoder aggregates multi-timestamp graph embeddings into the final timestamp $\mathbf{Z}_{T}$. (4) Structure and Attribute Reconstruction Decoder output reconstructed graph $\hat{\mathcal{E}}_{T}$ and node attributes $\hat{\mathbf{X}}_{T}$. (5) Anomaly scores $s_{v}$, calculated from the original and reconstructed graphs, identify and exclude the highest-scoring anomalous node from subsequent iterations.
4 Method
In this section, we present a graph-based method for detecting and mitigating hallucinations and errors in multi-agent collaboration by integrating information bottleneck theory with dual decoders.
4.1 Problem Formulation
Based on the temporal attributed graph framework, we aim to identify anomalous agents and their communications by analyzing both node attributes and structural patterns. This enables the removal of compromised agents and their connections, maintaining collaboration network integrity.
Given a temporal attributed graph sequence ${\{\mathcal{G}_{t}\}}_{t=1}^{T}$ where each $\mathcal{G}_{t}=(\mathcal{V}_{t},\mathcal{E}_{t},\mathbf{X}_{t})$, we formulate the anomaly detection objective as:
$\min_{f}\mathcal{L}(f)\quad\text{s.t.}$ $f:\mathcal{G}_{t}\rightarrow\mathbb{R}^{|\mathcal{V}_{t}|+|\mathcal{E}_{t}|},$ (1)
$\mathcal{V}_{t}^{*}=\{v\in\mathcal{V}_{t}\mid s_{v}>\tau\},$
where $s_{v}\in f(\mathcal{G}_{t})$, and $\mathcal{L}(f)$ denotes the loss function used to train the anomaly detection model. We further obtain the updated graph sequence $\mathcal{G}_{t}^{\prime}=(\mathcal{V}_{t}\setminus\mathcal{V}_{t}^{*},\mathcal{E}_{t}\setminus\mathcal{E}_{\mathcal{V}_{t}^{*}})$. To tackle this optimization problem, we focus on the carefully designed loss function $\mathcal{L}(f)$ that guides the anomaly detection model, and robust scoring mechanisms that compute node-level anomaly scores $s_{v}$. These scores enable the identification of anomalous nodes $\mathcal{V}_{t}^{*}$ through a reconstruction-based approach. The cleaned graph $\mathcal{G}_{t}^{\prime}$ is obtained by removing the identified anomalous elements.
Graph anomaly detection in multi-agent collaboration presents two critical challenges. First, traditional supervised approaches fail due to the inherent scarcity of labeled anomalies in dynamic temporal graphs. Second, the complex interplay between node attributed embeddings $\mathbf{X}_{t}$ and graph structure $\mathcal{E}_{t}$ creates a unique challenge, where anomalies can propagate through both attributes and topological dimensions simultaneously.
4.2 Encoder-Decoder Architecture
To overcome these limitations, we propose a novel unsupervised encoder-decoder architecture with dual decoders to capture the relationships between attribute and structural patterns. Our approach decomposes multi-agent system dynamics into four complementary components in Figure 3: (1) Attributed Graph Encoder that extends GCN capabilities [37] to capture structural and attribute correlations; (2) Time Information Encoder that adapts Transformer mechanisms [38] to integrate historical patterns into current timestep graph embeddings; (3) Attribute Reconstruction Decoder that reconstructs node attributes of the current timestep and preserves continuous feature fidelity; and (4) Structure Reconstruction Decoder that recovers the network topology of the current timestep and maintains discrete structural integrity. Therefore, this decomposition enables effective anomaly detection by independently reconstructing attribute space $\mathbf{X}_{t}$ and topological space $\mathcal{E}_{t}$, reducing the modal interference present in single-decoder approaches while allowing flexible weighting of different anomaly types across both spaces.
Attributed graph encoder. In multi-agent collaboration, node attributes and inter-agent connections are essential information carriers. At time step $t$, we employ GCN to learn node embeddings $\mathbf{Z}_{t}\in\mathbb{R}^{|\mathcal{V}_{t}|\times d}$ ($d$ is the embedding dimension) by aggregating information from neighboring nodes. Given node feature $\mathbf{X}_{t}\in\mathbb{R}^{|\mathcal{V}_{t}|\times k}$ ($k$ is feature dimension) and structure $\mathcal{E}_{t}$, GCN compresses node attributes and neighborhood information through the layer-wise propagation rule $\mathbf{H}^{(l+1)}=ReLU(\mathbf{D}_{e}^{-\frac{1}{2}}\mathbf{A}_{e}\mathbf{D}_{e}^{-\frac{1}{2}}\mathbf{H}^{(l)}\mathbf{W}^{(l)})$, where $\mathbf{H}^{(0)}=\mathbf{X}_{t}$. By stacking two GCN layers, nodes can perceive 2-hop neighborhood information while maintaining computational efficiency.
Time information encoder. In multi-agent systems, agents’ current decisions depend on historical interaction patterns. To capture these temporal dependencies, we employ a Transformer-based temporal encoder to process the sequence of graph embeddings $\{\mathbf{Z}_{1},\mathbf{Z}_{2},...,\mathbf{Z}_{T}\}$ obtained from the GCN encoder. Through the core self-attention mechanism $Attn(\mathbf{Q},\mathbf{K},\mathbf{V})=softmax(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}})\mathbf{V}$, where $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$ are query, key, and value matrices derived from the input sequence, the model dynamically weighs and aggregates information from previous interaction rounds, focusing on the most relevant historical patterns when generating the final representation $\mathbf{Z}_{T}$. This approach implicitly models temporal dependencies between nodes across different timesteps through cross-time self-attention, enabling the capture of how agent interactions evolve and influence each other over time.
Attribute reconstruction decoder. Node attributes in temporal graphs represent agent responses in multi-agent collaboration scenarios, encoding crucial semantic information. While structural information is discrete and binary, agent attributes are continuous and require specialized reconstruction mechanisms. We therefore implement a dedicated decoder for attribute reconstruction to preserve feature continuity and enable precise anomaly detection in attribute space.
Specifically, the attribute reconstruction decoder maps encoded latent representations $\mathbf{Z}_{T}$ to the reconstructed node attributes $\hat{\mathbf{X}}_{T}$. We compute the reconstruction error $\mathbf{R}_{\mathbf{X}}=\mathbf{X}_{T}-\hat{\mathbf{X}}_{T}$ to identify attribute-level anomalies, where larger errors indicate potential anomalous nodes. The reconstruction loss is measured using mean squared error: $\mathcal{L}_{\text{att}}=\frac{1}{|\mathcal{V}_{T}|}\sum_{i=1}^{|\mathcal{V}_{T}|}\|\bm{x}_{T,i}-\bm{\hat{x}}_{T,i}\|^{2}$.
Structure reconstruction decoder. Temporal graphs in agent collaboration systems face two interrelated challenges: anomalous connections that deviate from expected patterns, and redundant edges that introduce noise without contributing meaningful information. Our structure decoder addresses these challenges by analyzing binary adjacency matrices to identify both anomalous and redundant connections, thereby maintaining network efficiency and reliability.
The decoder reconstructs the network structure from latent representations $\mathbf{Z}_{T}$, modeling inter-agent communication patterns. We compute the structural reconstruction error $\mathbf{R}_{\mathcal{E}}=\mathcal{E}_{T}-\hat{\mathcal{E}}_{T}$, where a larger norm of $\mathbf{R}_{\mathcal{E}}(i,:)$ indicates a higher probability of structural anomalies for the $i$-th node. The reconstruction loss is measured by binary cross entropy: $\mathcal{L}_{\text{stru}}=-\frac{1}{{|\mathcal{V}_{T}|}^{2}}\sum_{(i,j)\in\mathcal{V}_{T}\times\mathcal{V}_{T}}[I_{(i,j)\in\mathcal{E}_{T}}\log(p_{ij})+I_{(i,j)\notin\mathcal{E}_{T}}\log(1-p_{ij})]$, where $I_{(\cdot)}$ denotes the indicator function and $p_{ij}$ means the predicted probability of edge $(i,j)$.
4.3 Graph Abstraction by Information Bottleneck
Dense connectivity in multi-agent frameworks creates structural complexity and information redundancy that obscures anomalous interactions. Our graph abstraction addresses this by: (1) capturing the relational nature of multi-agent systems, (2) enabling tracking of misinformation propagation paths, and (3) incorporating temporal dynamics to reveal evolutionary patterns that static models would miss. We apply Information Bottleneck Theory to balance information compression and preservation, reducing network complexity while maintaining essential temporal patterns. This temporal-aware abstraction forms the foundation for our analysis below.
Definition 4.1 (Temporal Graph Abstraction via Information Bottleneck).
Given a temporal graph sequence ${\{\mathcal{G}_{t}\}}_{t=1}^{T}$ where each $\mathcal{G}_{t}=(\mathcal{V}_{t},\mathcal{E}_{t},\mathbf{X}_{t})$ contains node set $\mathcal{V}_{t}$, edge set $\mathcal{E}_{t}$, and node features $\mathbf{X}_{t}\in\mathbb{R}^{|\mathcal{V}_{t}|\times d}$, the graph abstraction process seeks a compressed representation $\mathbf{Z}_{t}\in\mathbb{R}^{|\mathcal{V}_{t}|\times k}$ $(k<d)$ by minimizing:
$\mathcal{L}_{\text{GIB}}=I(\mathbf{X}_{t};\mathbf{Z}_{t})-\beta I(\mathbf{Z}_{t};\mathbf{Y}_{t})$ (2)
where $I(\mathbf{X}_{t};\mathbf{Z}_{t})$ controls the compression level, $I(\mathbf{Z}_{t};\mathbf{Y}_{t})$ ensures the retention of task-relevant information, and $\beta$ balances these objectives. This formulation enables the systematic reduction of graph complexity while maintaining essential temporal-structural patterns.
Remark 1.
In LLM multi-agent collaboration, $v_{t,i}\in\mathcal{V}_{t}$ represents the agent $m_{i}$ at time $t$, with features $\mathbf{X}_{t}$ encoding LLM’s response. Edge $e_{t,ij}\in\mathcal{E}_{t}$ represents inter-agent interactions. Target $\mathbf{Y}_{t}\in\mathcal{Y}$ denotes collaboration outcomes.
Theorem 4.2 (LLM Collaboration Information Bounds, Proof in Appendix A.2).
Under the GIB mechanism, the LLM multi-agent system provides the following theorem:
(1) Information Bottleneck: For any pair of collaborating agents, the information flow satisfies:
$I(\bm{x}_{t,i};\bm{x}_{t,j})\leq\eta I(\bm{x}_{t,i};\mathbf{Y}_{t})$
where $\eta$ is the controllable compression rate.
(2) Temporal Information Bottleneck: For the temporal evolution of agent representations, the mutual information between historical representations $\mathbf{Z}_{1:t-1}$ and current representations $\mathbf{Z}_{t}$ satisfies:
$I(\mathbf{Z}_{1:t-1};\mathbf{Z}_{t})\leq\mathbb{E}[\log P(\mathbf{Z}_{t}|\mathbf{Z}_{1:t-1})/Q(\mathbf{Z}_{t})]$
where $\mathbb{E}$ denotes the expectation, $P(\mathbf{Z}_{t}|\mathbf{Z}_{1:t-1})$ is the conditional probability and $Q(\mathbf{Z}_{t})$ is any prior distribution.
Remark 2.
This theorem establishes two key properties: (1) the information flow between agents is bounded by a controllable rate $\eta$, preventing error-amplifying cascade effects, and (2) the information flow between any agent’s historical and current states is bounded by a prior-guided constraint, ensuring smooth state transitions.
4.4 Incremental Training Paradigm
We adopt an incremental training paradigm that aligns with the multi-agent collaboration’s sequential nature. Unlike conventional unsupervised anomaly detection that randomly divides data, our approach leverages the temporal structure of interactions, using earlier timesteps to train for anomaly detection in later ones. For each new discussion round, our model processes merged interaction graphs from current and previous timesteps, with previously detected anomalous nodes removed. This ensures all collaborations undergo detection–an advantage over methods that assign some collaborations to training-only. By continuously fine-tuning across timesteps, our model builds progressively richer representations of normal behaviors while adapting to evolving interaction patterns, maintaining accuracy even as collaboration dynamics change.
Figure 4: A real case: co-existence of hallucination and agent-targeted errors.
The learning process is jointly guided by reconstruction errors and information bottleneck loss:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rec}}+\lambda\mathcal{L}_{\text{GIB}}$
where $\lambda$ balances the reconstruction fidelity and information compression objectives. The reconstruction loss $\mathcal{L}_{\text{rec}}$ combines attribute and structural components through:
$\mathcal{L}_{\text{rec}}=\alpha\mathcal{L}_{\text{att}}+(1-\alpha)\mathcal{L}_{\text{stru}}$
where $\alpha$ controls the relative importance between attribute and structural reconstruction.
Table 1: Accuracy (%) comparison of GPT-3.5-turbo, GPT-4o and Claude-3.5-sonnet under hallucination amplification or two types of error injection and propagation. Bold values represent the highest accuracy.
Method MMLU MATH FEVER
GPT-3.5-turbo GPT-4o Claude-3.5-sonnet GPT-3.5-turbo GPT-4o Claude-3.5-sonnet GPT-3.5-turbo GPT-4o Claude-3.5-sonnet
Hallucination Amplification
LLM Debate 54.5 80.1 77.3 34.6 52.3 57.3 30.6 33.3 33.1
DyLAN 56.3 83.3 78.3 40.8 76.4 75.6 32.3 37.4 37.2
SelfCheckGPT 55.1 82.2 77.5 7.4 51.3 42.7 3.3 3.6 33.6
GUARDIAN.s 56.2 86.4 80.2 49.3 76.6 75.6 34.1 40.4 38.5
GUARDIAN 57.2 84.9 82.3 56.2 78.5 79.2 34.5 41.8 39.2
Agent-targeted Error Injection and Propagation
LLM Debate 42.2 70.2 68.5 32.3 45.2 48.4 18.3 22.2 24.3
DyLAN 55.2 80.1 78.1 43.6 70.3 71.1 27.6 37.3 36.5
Challenger 31.8 45.2 42.3 36.3 49.3 52.1 17.2 20.8 21.3
Inspector 36.6 38.6 37.2 41.5 44.7 47.2 32.1 22.9 23.6
GUARDIAN.s 55.1 80.5 79.8 50.3 71.3 72.3 29.5 38.5 37.5
GUARDIAN 57.3 81.5 80.8 52.2 72.1 73.5 33.3 39.4 37.8
Communication-targeted Error Injection and Propagation
LLM Debate 37.2 78.2 75.7 31.1 51.1 52.4 30.3 23.5 25.6
DyLAN 52.6 81.2 78.5 41.3 76.3 74.2 34.1 36.5 37.9
Challenger 21.5 67.1 61.2 45.2 58.5 56.8 18.7 16.7 24.1
Inspector 33.5 77.3 73.6 46.5 60.2 62.4 31.6 24.5 29.4
GUARDIAN.s 56.6 82.5 78.2 54.2 77.3 73.8 35.1 38.1 38.5
GUARDIAN 60.1 83.7 79.1 53.9 78.4 75.2 35.3 38.6 39.3
5 Experiments
In this section, we present experimental settings and results for hallucination amplification and error injection and propagation scenarios. Our temporal graph model naturally aligns with multi-agent collaboration under the A2A protocol [12], with nodes representing agents and edges representing standardized communications.
5.1 Experimental Settings
Datasets. Our evaluation employs four benchmark datasets that span diverse domains and cognitive requirements. The benchmark datasets include MMLU [39], MATH [40], FEVER [41], and Biographies [1]. Specifically, these benchmarks test broad multi-subject knowledge (MMLU), mathematical reasoning (MATH), and factual verification against evidence (FEVER, Biographies). Following [11, 14], we randomly sample 100 questions from each dataset, with three independent testing iterations to ensure statistical robustness. We train a separate model for each dataset and apply the Incremental Training Paradigm within each dataset, focusing on in-distribution anomaly detection where the data distribution remains consistent across episodes.
Compared methods. We evaluate against three baseline categories: (1) standard multi-agent frameworks without defense (LLM Debate [1], DyLAN [11]); (2) hallucination detection methods (SelfCheckGPT [42]); and (3) error detection approaches (Challenger [30], Inspector [30]). This allows us to compare against both the inherent resilience of standard debate-based frameworks and dedicated defense strategies that employ self-consistency checks or deploy specialized inspector agents. LLM Debate serves as the fundamental framework, with several others building upon it. Details are in Appendix A.3. We present two variants: GUARDIAN.s performs anomaly detection using static graphs from the current time step, while GUARDIAN incorporates historical graph information.
Implementation details. We evaluate models in a zero-shot CoT [43] setting using both closed-sourced and open-sourced models (GPT-3.5-turbo [44], GPT-4o [45], Claude-3.5-sonnet [46], Llama3.1-8B [47]), with all agents treated without role differentiation. We primarily test with 4 agents, conducting additional experiments with 3-7 agents. Runtime efficiency is evaluated under communication-targeted attacks using consistent experimental settings. Detailed prompts are provided in Appendix A.4.
Evaluation metrics. We evaluate model accuracy, anomaly detection rate, False Discovery Rate (FDR), API calls, and runtime efficiency. For error injection and propagation, we simulate errors through agent-targeted attacks (randomly selecting 1 agent in the first round) and communication-targeted attacks (randomly interfering with multiple communication edges in intermediate rounds), simulating real-world malicious agents and communication interference, respectively. While the above tests utilize fully connected multi-agent topologies, we also conduct experiments with sparse communication graphs to validate our method’s effectiveness across different network structures, where each agent connects to only 25%, 50%, or 75% of agents in subsequent rounds. The detailed experimental setup is provided in Appendix A.5.
5.2 Experimental Results
Figure 5: API calls comparison across three scenarios: hallucination amplification and two types of error injection and propagation. Red values indicate the lowest number of API calls.
Hallucination amplification. In Table 1, we present the performance of LLM multi-agent collaboration models under hallucination amplification scenarios. Our model achieves an average improvement of $4.2\%$ over state-of-the-art baselines across multiple benchmarks. The gains are particularly pronounced on complex reasoning tasks, with a $7.1\%$ improvement on the MATH dataset over the best baseline across all backbone models, and reaching up to $15.4\%$ improvement when using GPT-3.5-turbo as the backbone. These results complement our theorems, where our GIB mechanism provides theoretical bounds on both the inter-agent information flow and temporal information propagation. This dual constraint ensures that both the spatial and temporal information flows remain bounded, while empirical results suggest these bounds contribute to reduced hallucination and effective collaboration. By treating LLMs as graph nodes, our approach remains model-agnostic and applicable across different architectures.
Table 2: Analysis of Detection Reliability: False Discovery Rate (FDR, %).
Safety Issue Datasets
MMLU MATH FEVER Biographies
Hallucination Amplification 26.67 13.11 17.86 15.69
Agent-targeted Error Injection and Propagation 22.22 8.32 20.53 13.23
Communication-targeted Error Injection and Propagation 30.67 18.42 28.57 19.65
Error injection and propagation. As shown in Table 1, GUARDIAN demonstrates superior defensive capabilities across attack scenarios. For agent-targeted attacks, our model achieves a $4.3\%$ improvement over the best baselines on MATH, and $8.6\%$ over GPT-3.5-turbo. In communication-targeted attacks, we achieve $3.6\%$ improvement over the best baselines on MMLU and MATH, and $7.5\%$ and $7.7\%$ respectively over GPT-3.5-turbo. These improvements stem from our encoder-decoder architecture that maps agent responses and communication patterns into a constrained latent space. This enables dual-level detection: reconstruction errors in node features signal agent-level attacks, while structural inconsistencies indicate communication-level threats. Our incremental training further enhances performance by processing interaction graphs sequentially, removing anomalous nodes while maintaining an evolving understanding of normal interactions and attack behaviors.
Our model’s anomaly detection rate results, detailed in Appendix A.6, average above 80% with a peak of 94.74%. We also analyzed potential failure modes where GUARDIAN may falsely remove correct agents. Table 2 presents False Discovery Rate (FDR) results with GPT-3.5-turbo across 4 agents, demonstrating consistently low rates across different scenarios, with an exceptionally low rate of 8.32% on the MATH dataset under agent-targeted attacks and remaining well below 20% for most scenarios. Additional analysis reveals that even when correct agents are mistakenly removed, the impact on overall system performance is minimal due to natural redundancy in multi-agent collaborations and GUARDIAN’s conservative strategy of removing only one agent per round, ensuring graceful degradation rather than catastrophic error amplification. Figure 4 shows a real case from the MATH dataset containing both hallucination and agent-targeted errors. In the first round, the four nodes fail to reach consensus, and GUARDIAN removes Node 1 as an anomaly due to its highest score. The remaining three nodes still fail to reach consensus in the second round, leading to the removal of Node 3. The final round between the two remaining nodes achieves consensus with the correct answer of 8. This process validates the effectiveness, as Node 1 was indeed a malicious agent and Node 3 exhibited hallucination.
Topology robustness. To further validate our approach’s robustness across different network topologies, Table 5.2 shows performance under sparse communication graphs in hallucination amplification scenarios, where agents connect to only 25%, 50%, or 75% of other agents in subsequent rounds. GUARDIAN consistently outperforms baselines across all sparsity levels, demonstrating effective performance beyond fully connected scenarios and validating our structure reconstruction decoder’s utility in non-trivial topologies.
Scalability analysis. To evaluate scalability, we varied the number of agents from 3 to 7 using GPT-3.5-turbo. As shown in Table 5.2, our framework maintains consistent performance across different configurations, demonstrating effective adaptation to multi-agent networks of varying sizes. This robust scalability stems from our temporal graph modeling approach, which efficiently captures agent interactions regardless of network size.
Running cost. As shown in Figure 5, our approach achieves optimal performance with the lowest API calls among all baselines. Our incremental node pruning strategy both removes anomalous nodes and naturally reduces redundant API queries, which is effective as agents tend to reach consensus in multi-round debates. Regarding runtime efficiency, Table 6 demonstrates that our method achieves the lowest average runtime per question on MMLU and FEVER datasets. On MATH, it incurs only marginal overhead (less than 5 seconds on average) compared to the baseline LLM Debate, while remaining significantly faster than other defense-enabled methods. The reduced runtime stems from our node deletion strategy, which decreases both API calls and inter-agent waiting time per round. Although anomaly detection is applied in each round, it introduces minimal computational overhead due to the small scale of agent communication graphs, contributing only a negligible fraction to the total runtime. Additional experimental results are provided in Appendix A.7,  A.8, and A.9.
Table 3: Accuracy (%) comparison under different connection sparsity under hallucination amplification on MATH dataset with GPT-3.5-turbo. Bold values represent the highest accuracy.
Method Sparsity
25% 50% 75%
LLM Debate 26.5 34.6 34.2
DyLAN 38.6 41.6 39.3
SelfCheckGPT 7.6 5.6 3.4
GUARDIAN 52.2 56.1 57.3
Table 4: Accuracy (%) of 3-7 agents on MATH dataset under hallucination amplification scenario, using GPT-3.5-turbo as the backbone.
Method Agent Number
3 4 5 6 7
LLM Debate 28.3 34.6 38.1 34.5 37.2
DyLAN 41.6 40.8 40.2 40.3 41.5
SelfCheckGPT 5.6 7.4 6.2 12.6 17.1
GUARDIAN.s 50.2 49.3 51.3 51.6 53.2
GUARDIAN 55.1 56.2 51.2 47.2 45.5
Table 5: Runtime cost (s) comparison under communication-targeted attacks with GPT-3.5-turbo. Bold values represent the lowest time cost.
Method MMLU MATH FEVER
LLM Debate 29.26 40.31 25.02
Challenger 26.15 56.23 27.5
Inspector 76.82 129.59 69.25
GUARDIAN 18.89 45.19 17.13
Figure 6: Parameter analysis of $\alpha$ and $\gamma$ for GUARDIAN accuracy.
5.3 Ablation Studies
We investigate two crucial parameters as shown in Figure 6: $\alpha$ controls the balance between structural and attribute reconstruction, while $\gamma$ regulates the compression-relevance trade-off in the information bottleneck (More details are presented in Appendix A.10). Using GPT-3.5-turbo with 4 agents, optimal performance is achieved when $\alpha\in[0.3,0.5]$ and $\gamma\in[0.001,0.01]$. These ranges suggest that a moderate $\alpha$ effectively balances attribute and structural information, preventing either aspect from dominating reconstruction. The optimal range of $\gamma$ ensures effective compression, preventing excessive information loss while avoiding retention of noise and redundant patterns in multi-agent interactions.
6 Conclusion
In this paper, we present GUARDIAN, a robust framework for addressing critical safety challenges in LLM-based multi-agent collaborations. By modeling interactions as temporal attributed graphs and leveraging an unsupervised encoder-decoder architecture, GUARDIAN enables effective anomaly detection without relying on modifying LLM models. Extensive experiments validate its state-of-the-art defensive performance with low computational cost.
Appendix A Technical Appendices and Supplementary Material
A.1 More Visualizations
We provide additional examples of graph-based visualization for LLM multi-agent collaboration in Figure 7 as a supplement to Figure 2, including hallucination amplification.
A.2 Proof
Proof.
(1) Information Bottleneck Guarantee: Consider two collaborating agents $m_{i}$ and $m_{j}$, where $\bm{z}_{t,j}$ represents the compressed information of agent $m_{j}$’s input $\bm{x}_{t,j}$, and $\mathbf{Y}_{t}$ is the target variable. By the data processing inequality, we have $I(\bm{x}_{t,i};\bm{x}_{t,j})\leq I(\bm{x}_{t,i};\bm{z}_{t,j})$. Then applying the Information Bottleneck principle yields $I(\bm{x}_{t,i};\bm{z}_{t,j})\leq\eta I(\bm{x}_{t,i};\mathbf{Y}_{t})$. Combining these inequalities directly leads to our desired result $I(\bm{x}_{t,i};\bm{x}_{t,j})\leq\eta I(\bm{x}_{t,i};\mathbf{Y}_{t})$.
(2) Temporal Information Bottleneck: By applying Lemma A.1 to variables $\mathbf{Z}_{1:t-1}$ and $\mathbf{Z}_{t}$, we complete the proof [48, 49]). ∎
Lemma A.1 (Mutual Information Upper Bound in VIB).
Given any two variables X and Y, we have the variational upper bound of I(X;Y):
$I(X;Y)$ $=\mathbb{E}_{P(X,Y)}[\log\frac{P(Y|X)}{P(Y)}]$
$=\mathbb{E}_{P(X,Y)}[\log\frac{P(Y|X)Q(Y)}{P(Y)Q(Y)}]$
$=\mathbb{E}_{P(X,Y)}[\log\frac{P(Y|X)}{Q(Y)}]-\underbrace{KL[P(Y)||Q(Y)]}_{\text{non-negative}}$
$\leq\mathbb{E}_{P(X,Y)}[\log\frac{P(Y|X)}{Q(Y)}]$
A.3 More Information about Baselines
We compare our approach with the following baselines, which can be categorized into three groups: fundamental multi-agent frameworks, hallucination detection methods, and error detection approaches.
Fundamental LLM Multi-agent Frameworks
  • LLM Multi-agent Debate [1], first proposed in this research, enhances model performance through structured discourse between multiple instances, yielding improved factual accuracy and reasoning through iterative consensus building.
  • DyLAN [11] is a framework that builds a dynamic multi-agent feed-forward network structure, automatically selecting and optimizing agent teams during inference, and improving LLM multi-agent collaboration performance and efficiency through consensus-based early stopping.
Figure 7: Hallucination amplification visualization.
Hallucination Detection Methods
  • SelfCheckGPT [42] is a classic method for detecting hallucinations in standalone LLMs by comparing a model’s multiple responses to evaluate their consistency. Specifically, it uses natural language prompts to perform consistency checks between a main response and several randomly generated sample responses. Through prompt templates, it guides LLM to determine whether sentences are supported (answering Yes or No), quantifying results into consistency scores. Lower scores indicate sentences are more likely to be factual; higher scores suggest potential hallucinations. We integrate this self-checking mechanism into LLM multi-agent collaboration as a baseline approach. When integrated into multi-agent collaboration scenarios, each agent reviews answers from other agents (excluding the one with the highest hallucination score as determined by SelfCheckGPT), and rethinks and responds based on these references. The system performs hallucination detection on each agent’s response to ensure reliability. Finally, the system aggregates all answers that pass the hallucination detection threshold and employs majority voting to reach the final judgment.
Error Detection Methods
  • Challenger [30] enhances existing agents through configuration file modifications, enabling them to question and verify other agents’ outputs. Each agent evaluates received messages for potential errors before task execution, responding with "safe" or "unsafe" accordingly.
  • Inspector [30] introduces a specialized oversight agent that monitors inter-agent communications, functioning as an independent validator to intercept and analyze message exchanges for potential errors and inconsistencies.
A.4 Prompt Templates
Figure 8 illustrates all prompt templates for different safety scenarios. Figure 9 shows the role-specific prompts. While roles are not differentiated within each task, they are distinguished across different tasks.
Figure 8: Prompt templates for different safety scenarios.
Figure 9: Role-specific prompt templates.
Table 6: Anomaly Detection Rate (%) on different datasets.
Safety Issue Datasets
MMLU MATH FEVER Biographies
Hallucination Amplification 71.27 82.77 91.52 66.92
Agent-targeted Error Injection and Propagation 83.99 85.94 92.63 72.31
Communication-targeted Error Injection and Propagation 72.21 76.73 94.74 70.65
A.5 More Information about Setup
The benchmark datasets include: MMLU [39], which tests broad knowledge across 57 subjects from elementary to professional levels through multiple-choice questions; MATH [40], which challenges mathematical reasoning with 12,500 competition problems and their step-by-step solutions; FEVER [41], which consists of 185,445 Wikipedia-derived claims labeled as Supported, Refuted, or NotEnoughInfo, along with supporting evidence for verifiable claims; and Biographies [1], which contains ground truth data of 524 computer scientists to evaluate models’ accuracy in biographical information generation. FEVER and Biographies datasets are particularly prone to factual inaccuracies [50].
For evaluation metrics, we report model accuracy, anomaly detection rate, False Discovery Rate (FDR), the number of API calls, and runtime efficiency. Model accuracy represents the correctness of final answers produced by the multi-agent system, serving as our primary evaluation metric. The anomaly detection rate measures the system’s ability to identify anomalous patterns during inter-agent interactions [51]. Specifically, we employ a Time-decaying Weighted Anomaly Detection Rate that assigns greater importance to anomalies detected in earlier interaction rounds. This weighting mechanism is implemented through either exponential or linear decay functions, where anomalies in earlier rounds receive higher weights. This approach is particularly valuable for our scenario, as it emphasizes the system’s capability to identify hallucination amplification or anomaly propagation at early stages. For each round, if the detected anomalous node is indeed anomalous (true positive), we assign a score of 1; otherwise, 0. The final anomaly detection rate is calculated as the weighted average across all rounds. The False Discovery Rate (FDR) directly addresses a critical practical question: “When GUARDIAN flags a node as anomalous and intervenes, what is the probability that this is a false alarm?” Technically, FDR is calculated as the ratio of false positives (nodes incorrectly flagged as anomalous) to the total number of detections, measuring the precision and trustworthiness of our system’s interventions. API calls serve as a measure of communication efficiency between agents. We use API calls rather than token consumption alone, as token usage shows high variability across different tasks and prompting strategies and thus cannot reliably reflect communication efficiency. Runtime efficiency is measured as the average time per question in seconds, evaluating the computational overhead of our approach compared to baseline methods.
Table 7: Accuracy (%) and API calls comparison on GPT-3.5-turbo with different numbers of agents. Bold values represent the highest accuracy and lowest API calls, respectively.
Dataset Agent Metric LLM Debate DyLAN SelfCheckGPT GUARDIAN.s GUARDIAN
acc(%) 54.2 56.1 56.3 56.2 56.6
3 API 6.00 4.77 12.00 3.82 4.02
acc(%) 54.5 56.3 55.1 56.2 57.2
4 API 8.00 5.27 16.00 4.52 4.48
acc(%) 51.3 57.3 54.2 57.2 60.3
5 API 10.00 5.26 20.00 5.47 5.74
acc(%) 53.5 56.3 54.8 55.8 57.2
6 API 12.00 6.77 24.00 6.04 6.83
acc(%) 53.4 55.9 56.3 56.2 57.3
MMLU 7 API 14.00 6.31 28.00 6.02 6.20
acc(%) 28.3 41.6 5.6 50.2 55.1
3 API 6.00 8.46 12.00 4.92 4.95
acc(%) 34.6 40.8 7.4 49.3 56.2
4 API 8.00 3.18 16.00 6.22 5.89
acc(%) 38.1 40.2 6.2 51.3 51.2
5 API 10.00 4.76 20.00 7.34 6.29
acc(%) 34.5 40.3 12.6 51.6 47.2
6 API 12.00 5.54 24.00 14.58 8.61
acc(%) 37.2 41.5 17.1 53.2 45.5
MATH 7 API 14.00 5.29 28.00 15.33 6.80
acc(%) 51.4 92.5 16.8 95.8 95.6
3 API 6.00 12.00 12.00 6.00 6.00
acc(%) 52.5 93.1 19.4 95.3 95.6
4 API 8.00 16.00 16.00 9.00 9.00
acc(%) 52.5 92.1 25.2 95.1 94.9
5 API 10.00 19.80 20.00 14.20 10.78
acc(%) 50.3 93.2 36.3 96.8 95.4
6 API 12.00 23.50 24.00 20.60 11.30
acc(%) 51.1 93.5 32.5 96.1 97.2
Biographies 7 API 14.00 27.80 28.00 28.36 11.56
acc(%) 28.2 34.5 3.5 35.1 34.6
3 API 6.00 5.16 12.00 3.93 4.21
acc(%) 30.6 32.3 3.3 34.1 34.5
4 API 8.00 3.95 16.00 4.43 4.11
acc(%) 31.2 34.2 3.1 37.3 36.3
5 API 10.00 6.03 20.00 7.02 5.03
acc(%) 30.8 31.3 3.9 35.2 32.6
6 API 12.00 6.23 24.00 9.83 5.78
acc(%) 30.3 32.6 5.4 33.5 34.1
FEVER 7 API 14.00 7.53 28.00 8.53 7.02
A.6 Additional Experiments on Anomaly Detection Rate
Table 6 presents results of anomaly detection rate with GPT-3.5-turbo across 4 agents.
A.7 Additional Experiments on Hallucination Amplification
Table 7 presents results with GPT-3.5-turbo across 3-7 agents. Table A.7 shows results using Llama3.1 8B with 4 agents. Table A.7 shows results on the Biographies dataset. GUARDIAN achieves highest robustness in most scenarios, with GUARDIAN.s also showing strong performance.
Table 8: Accuracy (%) and API calls comparison on Llama3.1-8B. Bold values represent the highest accuracy and lowest API calls, respectively.
Method Metric MMLU MATH FEVER
LLM Debate Acc.(%) 48.4 24.2 27.5
API 12.00 12.00 12.00
DyLAN Acc.(%) 50.1 27.8 39.8
API 4.74 6.89 6.36
GUARDIAN.s Acc.(%) 57.2 28.5 40.2
API 4.81 6.25 4.47
GUARDIAN Acc.(%) 52.5 30.2 42.6
API 4.56 6.58 4.45
Table 9: Accuracy (%) and API calls comparison between GPT-3.5-turbo, GPT-4o, and Claude-3.5-sonnet on the Biographies dataset. Bold values represent the highest accuracy and lowest API calls, respectively.
Method Metric Biographies
GPT-3.5-turbo GPT-4o Claude-3.5-sonnet
LLM Debate Acc.(%) 52.5 94.6 94.5
API 8.00 8.00 8.00
DyLAN Acc.(%) 93.1 96.5 96.3
API 12.00 12.00 12.00
SelfCheckGPT Acc.(%) 19.4 91.3 88.3
API 16.00 16.00 16.00
GUARDIAN.s Acc.(%) 95.3 97.2 96.5
API 9.00 9.00 9.00
GUARDIAN Acc.(%) 95.6 97.5 96.9
API 9.00 9.00 9.00
Table 10 presents the API call statistics for hallucination amplification.
Table 10: API calls comparison between GPT-3.5-turbo, GPT-4o and Claude-3.5-sonnet. Bold values represent the lowest API calls.
Method MMLU MATH FEVER
GPT-3.5-turbo GPT-4o Claude-3.5-sonnet GPT-3.5-turbo GPT-4o Claude-3.5-sonnet GPT-3.5-turbo GPT-4o Claude-3.5-sonnet
LLM Debate 8.00 8.00 8.00 8.00 8.00 8.00 8.00 8.00 8.00
DyLAN 5.27 3.15 3.30 3.18 4.24 4.20 3.95 3.37 3.22
SelfCheckGPT 16.00 16.00 16.00 16.00 16.00 16.00 16.00 16.00 16.00
GUARDIAN.s 4.52 3.17 3.27 6.22 4.09 4.46 4.43 3.21 3.31
GUARDIAN 4.48 3.25 3.23 5.89 3.81 4.39 4.11 3.26 3.16
Table 11: API calls comparison between GPT-3.5-turbo, GPT-4o and Claude-3.5-sonnet under agent-targeted attacks. Bold values represent the lowest API calls.
Method MMLU MATH FEVER
GPT-3.5-turbo GPT-4o Claude-3.5-sonnet GPT-3.5-turbo GPT-4o Claude-3.5-sonnet GPT-3.5-turbo GPT-4o Claude-3.5-sonnet
LLM Debate 8.00 8.00 8.00 8.00 8.00 8.00 8.00 8.00 8.00
DyLAN 4.36 4.19 4.56 5.10 4.50 4.95 4.90 4.16 3.98
Challenger 8.00 8.00 8.00 8.00 8.00 8.00 8.00 8.00 8.00
Inspector 16.00 16.00 16.00 16.00 16.00 16.00 16.00 16.00 16.00
GUARDIAN.s 5.30 4.35 4.78 6.40 4.85 4.86 5.07 4.06 4.02
GUARDIAN 5.07 4.17 4.51 6.45 4.44 4.36 5.09 3.98 3.74
Table 12: API calls comparison between GPT-3.5-turbo, GPT-4o and Claude-3.5-sonnet under communication-targeted attacks. Bold values represent the lowest API calls.
Method MMLU MATH FEVER
GPT-3.5-turbo GPT-4o Claude-3.5-sonnet GPT-3.5-turbo GPT-4o Claude-3.5-sonnet GPT-3.5-turbo GPT-4o Claude-3.5-sonnet
LLM Debate 12.00 12.00 12.00 12.00 12.00 12.00 12.00 12.00 12.00
DyLAN 4.16 3.37 3.75 4.80 3.60 4.36 3.95 3.85 3.97
Challenger 12.00 12.00 12.00 12.00 12.00 12.00 12.00 12.00 12.00
Inspector 24.00 24.00 24.00 24.00 24.00 24.00 24.00 24.00 24.00
GUARDIAN.s 4.02 3.56 3.66 5.85 3.79 4.02 4.23 3.51 3.63
GUARDIAN 4.30 3.36 3.30 6.02 4.18 3.98 3.95 3.27 3.20
A.8 Additional Experiments on Error Injection and Propagation
The communication-targeted attacks are implemented in debate scenarios with a minimum of 3 rounds. The perturbation is introduced between rounds 1 and 2, targeting all incoming communication edges of a selected agent in round 2.
Table 11 and  12 present the API call statistics for error injection and propagation. Table A.8 and A.8 show results on the Biographies dataset.
Table 13: Accuracy (%) and API calls comparison between GPT-3.5-turbo, GPT-4o, and Claude-3.5-sonnet under agent-targeted attacks on the Biographies dataset. Bold values represent the highest accuracy and lowest API calls, respectively.
Method Metric Biographies
GPT-3.5-turbo GPT-4o Claude-3.5-sonnet
LLM Debate Acc.(%) 40.5 91.3 85.3
API 8.00 8.00 8.00
DyLAN Acc.(%) 82.1 93.2 86.2
API 12.00 12.00 12.00
Challenger Acc.(%) 40.5 87.7 80.7
API 8.00 8.00 8.00
Inspector Acc.(%) 39.4 90.3 86.3
API 16.00 16.00 16.00
GUARDIAN.s Acc.(%) 88.1 93.1 87.1
API 9.00 9.00 9.00
GUARDIAN Acc.(%) 88.4 93.7 87.6
API 9.00 9.00 9.00
Table 14: Accuracy (%) and API calls comparison between GPT-3.5-turbo, GPT-4o, and Claude-3.5-sonnet under communication-targeted attacks on the Biographies dataset. Bold values represent the highest accuracy and lowest API calls, respectively.
Method Metric Biographies
GPT-3.5-turbo GPT-4o Claude-3.5-sonnet
LLM Debate Acc.(%) 54.8 55.6 60.3
API 12.00 12.00 12.00
DyLAN Acc.(%) 94.1 94.2 87.5
API 12.00 12.00 12.00
Challenger Acc.(%) 58.9 57.1 58.6
API 12.00 12.00 12.00
Inspector Acc.(%) 43.2 56.4 64.6
API 24.00 24.00 24.00
GUARDIAN.s Acc.(%) 95.7 96.2 87.3
API 9.00 9.00 9.00
GUARDIAN Acc.(%) 94.9 96.5 89.8
API 9.00 9.00 9.00
A.9 More Real Cases
Addressing the safety issues presented in Appendix A.1 and Figure 2(b), we demonstrate how GUARDIAN handles these challenges, with results shown as follows: Figure 10 addresses hallucination amplification (Figure 7), Figure 11 resolves agent-targeted error injection and propagation, and Figure 12 mitigates communication-targeted error injection and propagation (Figure 2(b)).
Figure 10: A real case of hallucination amplification and how GUARDIAN addresses it.
Figure 11: A real case of agent-targeted error injection and propagation and how GUARDIAN addresses it.
Figure 12: A real case of communication-targeted error injection and propagation and how GUARDIAN addresses it.
A.10 Details about Ablation Studies
The hyperparameter $\gamma$ is derived from the trade-off between reconstruction and information compression. Our total loss function is $L_{\text{total}}=L_{\text{rec}}+\lambda(I(\mathbf{X};\mathbf{Z})-\beta I(\mathbf{Z};\mathbf{Y}))$. Using the approximation $L_{\text{rec}}\approx-I(\mathbf{Z};\mathbf{Y})$ from prior work [48], the objective can be rewritten as:
$L_{\text{total}}\approx\lambda I(\mathbf{X};\mathbf{Z})+(1+\lambda\beta)L_{\text{rec}}$
By normalizing the total loss, we can express the relationship as:
$\frac{L_{\text{total}}}{1+\lambda\beta}\approx\gamma I(\mathbf{X};\mathbf{Z})+L_{\text{rec}},\quad\text{where}\quad\gamma=\frac{\lambda}{1+\lambda\beta}$
Here, $\gamma$ directly controls the weight of the information compression term relative to the reconstruction loss.